# Self-Supervised Millimeter Wave Indoor Localization using Tiny Neural Networks

### v1

@@ V-B Performance evaluation for entire dataset
We now evaluate the performance of our tiny NN. As NN outperform geometry-based localization schemes [13], we evaluate our model against JADE. For this, we exploit the entire set of receiver locations across the room. Specifically, we initially run JADE on the entire dataset of $\sim$2 600 measurements to obtain the corresponding RX location estimates. Even when working on datasets with many measurements, some of JADE’s location estimates can be largely off. To avoid that such errors negatively affect the NN training process, we compute the centroid of all location estimates, fit a Gaussian distribution to the location estimates, and remove the 5% of all estimates that are farthest from the centroid. We then form a test set with 434 location labels and their corresponding ADoA features, and train the NN over the remaining points.
@@
We illustrate the performance of our tiny NN through the cumulative distribution function (CDF) of the RX location error in Fig. 5, where we also compare our self-supervised approach with the case the NN model is trained with ground-truth location labels. The NN model used has (40, 36, 36, 18, 2)403636182(40,\ 36,\ 36,\ 18,\ 2)( 40 , 36 , 36 , 18 , 2 ) neurons respectively in each layer with the optimal hyperparameters $\kappa=0.9$, $p=0.2$, $r=0.004$, and $b=75\%$. We observe that our tiny self-supervised NN performs similar to JADE. Both schemes achieve sub-meter localization accuracy in about $74\%$ of the cases. When trained with true RX labels, the NN outperforms JADE, with sub-meter accuracy in about $88\%$ of the cases. We emphasize that training the NN with ground-truth location labels is used here only to verify that the NN learns to accurately map ADoA measurements to the corresponding RX location. In practice, ground-truth locations would require a burdensome training dataset collection process, making our self-supervised approach much more convenient. While JADE and our NN perform comparably in Fig. 5, JADE’s complexity increases as new measurements flow in, whereas the NN only requires a limited set of matrix multiplications and vector additions computed on a small-sized input set, once the training is complete.
Figure 5: CDF of the location estimation error for the entire dataset.
Figure 6: CDF of the location estimation errors for our proposed 3-hidden layer (HL) NN against the 2-hidden layer NN.
We now turn to assessing the impact of the number of hidden layers (HL) in the NN, as this parameter affects the computational complexity of the NN. We thus compare the model proposed in Section III-D, consisting of 3 hidden layers, with the NN model proposed in [13], consisting of 2 hidden layers. We train and test both NNs in self-supervised manner.
@@
The location error CDFs in Fig. 6 show that our 3-hidden layer NN model performs similar to JADE and outperforms the 2-hidden layer version, which attains sub-meter errors only in $65\%$ of the cases. In this challenging environment with a non-uniform distribution of both the RX locations across different regions and the number of ADoA measurements at different RX locations, adding a hidden layer improves the localization accuracy by learning a more complex relationship between ADoA values and the location of the RX. This contrasts the 2-layer NN in [13], which worked well only with uniform datasets and regular room shapes.
Although adding hidden layers may lead to more accurate localization, it would also increase the computation of NN training and inference. As our primary goal is to learn the smallest NN model that can match the performance of JADE, we limit our model search to just 3 hidden layers.
@@
Figure 7: Localization error for all regions with different number of transmitters available. (a) Location-wise performance of self-supervised tiny NNs at the test locations (b) Statistical dispersion of the location error for regions having different numbers of available transmitters.
@@
To better understand the performance of our NN, we examine how localization errors distribute over different locations across the room. Fig. 7a codes localization error at each test location via a colored dot, where blue hues represent low or acceptable errors. We observe localization errors of <1 m at the majority of the test locations, mostly in areas 2, 4, 5, and 6. There are two reasons behind this result. First, the RX moving in these areas receives signals from 3 or 4 TXs. This yields a larger number of clusterized MPCs, thus a larger number of input ADoA values. The greater the number of ADoAs per location, the more accurately JADE and the NN can localize the RX. Secondly, these areas include a larger number of RX measurements available compared to areas 1 and 3, and constitute a larger proportion of the training data. Thus, the model can estimate the location of the RX with better accuracy. On the contrary, large estimation errors in areas 1 and 3 tend to be larger. This is because the RX records measurements only from two transmitters, and the set of measurement locations is sparser compared to other areas. However, many locations in these areas still yield sub-meter errors, implying that both JADE and our self-supervised NN infer the relationship between ADoAs and RX locations even by observing a limited number of examples.
@@ V-C Performance evaluation over different areas
Notice that the performance of both JADE and the NN are better in areas 2, 3 and 4 as well as area 5, and more erroneous in area 6. This is a consequence of resorting to a self-supervised approach, which inherently achieves the best performance only when the bootstrapping algorithm is also accurate throughout a given area. In our case, JADE performed remarkably good in areas 2, 3 and 4 as well as area 5. Here, the NN is trained with low-error location labels, and therefore performs well. Conversely, JADE’s location labels are affected by larger errors in area 6. This affects the training of the NN, which achieves sub-meter errors in 75% of the cases, instead of, e.g., 85% for area 5.
Compare now against the performance of the NN trained with ground-truth location labels: we observe that the overall accuracy loss when the bootstrapping algorithm is accurate (such as in areas 2, 3 and 4, or in area 5) brings the proportion of locations with sub-meter errors from about $\approx$90to $\approx$85%. If the bootstrapping algorithm is less accurate (e.g., like JADE working on data from area 6), the proportion decreases to about 75%. Even though area 6 provides the largest number of dataset points, compounding training data from all areas increases the localization performance of the NN with respect to training on data from area 6 alone.
@@
V-D Simulation study of localization error distribution
@@
Fig. 9b shows the CDF of the location error as estimated using JADE and the self-supervised NN. We observe that the NN performs as good as JADE with sub-meter location accuracy in about 92% of the cases, with slightly better performance for errors less than $50$ cm. For comparison, 98% of the estimates yield an error of <1 m when the NN is trained with true locations. These results are significantly good for two main reasons: because JADE can work with a comparatively large measurement set, and because each measurement location is illuminated by multiple MPCs. Bootstrapping with accurate location labels from JADE thus improves the performance of the NN.
@@
Fig. 9c illustrates the localization error heat map in the H-shaped room, for all test locations, as estimated using the self-supervised NN. Blue hues represent low estimation errors (<1 m), light blue to green hues represent errors between 1 m and 2 m, whereas yellow to red hues convey errors between 3 m and 6 m. The results confirm that the localization error is <1 m throughout the room, except for slightly higher errors in the bottom-right section. A similar situation occurs near the top-right corner of the middle section. The main reason is that the training dataset is not collected at uniformly distributed locations, and may thus be sparser in some areas, which also explains why the heat map is not symmetrical.
@@
In general our performance evaluation shows that a tiny self-supervised NN can accurately localize a user in a complex indoor environment in the presence of two key factors: $i)$ richness of input features, coming either from LoS paths, NLoS paths, or both; and $ii)$ richness of training data, so that the NN observes sufficiently many error-prone samples.
@@
VI Conclusions
We presented a self-supervised approach to train a tiny NN to localize a mmWave device using ADoA measurements in an indoor environment. This relieves the burdensome task of collecting training data by harvesting location labels from a bootstrapping localization algorithm. We evaluated the performance of our scheme via channel measurements from a NIST mmWave channel sounder. The channel measurements were processed using the SAGE algorithm to obtain all the MPCs at the receiver. To identify dominant MPCs reflected off different surfaces, we proposed a recursive clustering algorithm based on DBSCAN. The azimuthal AoAs associated with the cluster centroids were used to compute ADoAs. Our performance evaluation shows that tiny NNs can accurately estimate the location of the receiver when trained with true locations. While the error inevitably increases in the self-supervised case, our scheme still achieves sub-meter error in about 75% of the cases, with a much lower computational complexity than the bootstrapping algorithm. Additional simulation results confirm that the key elements that enable high accuracy are richness of illumination from multiple MPCs and training data.

### v2

@@ V-B Performance evaluation for entire dataset
We now evaluate the performance of our tiny NN. We evaluate our model against JADE and other geometry-based localization schemes. For this, we exploit the entire set of receiver locations across the room. Specifically, we initially run JADE on the entire dataset of $\sim$2 600 measurements to obtain the corresponding RX location estimates. Even when working on datasets with many measurements, some of JADE’s location estimates can be largely off. To avoid that such errors negatively affect the NN training process, we compute the centroid of all location estimates, fit a Gaussian distribution to the location estimates, and remove the 5% of all estimates that are farthest from the centroid. We then form a test set with 434 location labels and their corresponding ADoA features, and train the NN over the remaining points.
@@
We illustrate the performance of our tiny NN through the cumulative distribution function (CDF) of the RX location error in Fig. 5, where we also compare our algorithm-supervised approach (referred in the figures as AS-TNN) with the case when the NN model is trained with ground-truth location labels (referred as TNN). Moreover, we also compare against JADE, the 2-hidden layer NN model proposed in [13], and the geometric scheme ADOA [18]. Our proposed NN model has (40, 36, 36, 18, 2)403636182(40,\ 36,\ 36,\ 18,\ 2)( 40 , 36 , 36 , 18 , 2 ) neurons respectively in each layer with the optimal hyperparameters $\kappa=0.9$, $p=0.2$, $r=0.004$, and $b=75\%$. We observe that our tiny algorithm-supervised NN performs similar to JADE. Both schemes achieve sub-meter localization accuracy in about $74\%$ of the cases, with mean and median localization errors of about 86 cm and $\sim$61 cm, respectively. On the other hand, the previous 2-hidden layer version of the tiny NN achieves sub-meter errors in 65% of the estimates, with the corresponding mean and the median errors being 98 cm and $\sim$74 cm.
@@
Our test environment entails a non-uniform distribution of RX locations across the room and a different number of ADoA measurements in each room area. In this case, adding a hidden layer improves the localization accuracy by learning a more complex relationship between ADoA values and the location of the RX. This contrasts the 2-layer NN in [13], which worked well only with uniform datasets and regular room shapes. We observe that our proposed models as well as the zero-knowledge algorithm JADE outperform the geometric algorithm ADoA, that achieves sub-meter errors only in 40% of the cases.
When trained with true RX labels, the NN outperforms JADE, with sub-meter accuracy in about $88\%$ of the cases, and a lower mean and median errors of 60 cm and $\sim$45 cm. We emphasize that training the NN with ground-truth location labels is used here only to verify that the NN learns to accurately map ADoA measurements to the corresponding RX location. In practice, ground-truth locations would require a burdensome training dataset collection process, making our algorithm-supervised approach much more convenient. While JADE and our NN perform comparably in Fig. 5, JADE’s complexity increases as new measurements flow in [11], whereas the NN only requires a limited set of matrix multiplications and vector additions computed on a small-sized input set, once the training is complete.
Figure 5: CDF of the location estimation errors for our proposed 3-hidden layer (HL) NN (referred to as AS-TNN) against the previous 2-hidden layer model and other state-of-the-art algorithms. Note that our tiny NN trained with ground truth is referred to as TNN.
@@
Figure 6: Localization error for all regions with different number of transmitters available. (a) Location-wise performance of algorithm-supervised tiny NNs at the test locations (b) Statistical dispersion of the location error for regions having different numbers of available transmitters.
@@
To better understand the performance of our NN, we examine how localization errors distribute over different locations across the room. Fig. 6a codes localization error at each test location via a colored dot, where blue hues represent low or acceptable errors. We observe localization errors of <1 m at the majority of the test locations, mostly in areas 2, 4, 5, and 6. There are two reasons behind this result. First, the RX moving in these areas receives signals from 3 or 4 TXs. This yields a larger number of clusterized MPCs, thus a larger number of input ADoA values. The greater the number of ADoAs per location, the more accurately JADE and the NN can localize the RX. Secondly, these areas include a larger number of RX measurements available compared to areas 1 and 3, and constitute a larger proportion of the training data. Thus, the model can estimate the location of the RX with better accuracy. On the contrary, large estimation errors in areas 1 and 3 tend to be larger. This is because the RX records measurements only from two transmitters, and the set of measurement locations is sparser compared to other areas. However, many locations in these areas still yield sub-meter errors, implying that both JADE and our algorithm-supervised NN infer the relationship between ADoAs and RX locations even by observing a limited number of examples.
@@ V-C Performance evaluation over different areas
Notice that the performance of both JADE and the NN are better in areas 2, 3 and 4 as well as area 5, and more erroneous in area 6. This is a consequence of resorting to an algorithm-supervised approach, which inherently achieves the best performance only when the bootstrapping algorithm is also accurate throughout a given area. In our case, JADE performed remarkably good in areas 2, 3 and 4 as well as area 5. Here, the NN is trained with low-error location labels, and therefore performs well. Conversely, JADE’s location labels are affected by larger errors in area 6. This affects the training of the NN, which achieves sub-meter errors in 75% of the cases, instead of, e.g., 85% for area 5.
Compare now against the performance of the NN trained with ground-truth location labels: we observe that the overall accuracy loss when the bootstrapping algorithm is accurate (such as in areas 2, 3 and 4, or in area 5) brings the proportion of locations with sub-meter errors from about $\approx$90to $\approx$85%. If the bootstrapping algorithm is less accurate (e.g., like JADE working on data from area 6), the proportion decreases to about 75%. Even thxough area 6 provides the largest number of dataset points, compounding training data from all areas increases the localization performance of the NN with respect to training on data from area 6 alone.
@@
V-D Impact of number of hidden layers and clustering of MPCs
We now turn to assessing the impact of the number of hidden layers in the NN, as this parameter affects the computational complexity. Fig. 8 presents the statistical dispersion of localization errors for varying number of hidden layers in the tiny NN model. Specifically, we vary from 2 hidden layers (as proposed in [13]) to 6. We observe a decrease in the median error from $\sim$76 cm to $\sim$60 cm as we move from a 2-HL model to a 3-HL model. As more layers are added to the model, the whiskers of the boxplot span almost the same interval, implying a similar performance as a 3-HL model. We also note that, as the number of layers increases, the NN will be able to represent a more complex regression function, but at a cost of learning and tuning more parameters. Moreover, adding hidden layers may lead to more accurate localization, it would also increase the computational complexity of NN training and inference. As our primary goal is to learn the smallest NN model that can match the performance of JADE, we limit our search to a 3-layer model.
Figure 8: Statistical dispersion of localization errors for varying number of hidden layers in the tiny NN. Here, AS-TNN - kHL denotes our algorithm-supervised tiny NN with k hidden layers.
Figure 9: CDF of the location errors when using the dominant MPCs obtained from the regular DBSCAN and the proposed recursive DBSCAN (here, Rec-DBSCAN).
To understand the impact of clustering the MPCs on the localization performance, we propose to feed JADE and our NN model with ADoAs obtained after both regular and our proposed recursive DBSCAN procedure. Fig. 9 presents the CDF of the localization errors incurred by JADE as well as our algorithm-supervised tiny NN model when using ADoAs obtained from the two procedures. While the NN model is able to match the performance of the bootstrapping localization algorithm in both the cases, we can clearly observe an increase in the sub-meter localization errors from 61% to 74% when recursively clustering the MPCs. This is because recursive DBSCAN finely breaks down large clusters, resulting in rich set of ADoA measurements corresponding to the dominant MPCs reflected off the walls. This significantly improves JADE’s performance, and subsequently that of the TNN.
V-E Complexity analysis of the tiny NN
We now evaluate the complexity of our NN model using the following two metrics:
[…]
Mean training and testing times. We perform preliminary experiments to evaluate the mean training and inference times of our NN model. For this, we used a Thinkpad E14 laptop, with an AMD Ryzen 7 4000 CPU and 24 GBytes of RAM, running in battery-saver mode. Our NN model takes an average of $\sim$165 s to train, and the mean inference time is about 1.22 ms. Running JADE on the same machine takes about 1 400 s (about 23 mins) to obtain the training labels. This is because JADE solves a large grid search problem having $2^{16}$ points, followed by computationally intensive iterative solution of two inter-dependent MMSE problems. Thus, from a systems perspective, we propose to offload JADE and the NN training process to an edge server.
We remark that as the indoor scenario becomes more complex (i.e., with more obstacles, complex-shaped rooms, additional APs, etc.), the number of MPCs between the TX and the RX will increase. To accurately localize the client, our bootstrapping localization algorithm will have to incorporate more measurements for each location, and process them altogether with a higher degree of computational complexity. As the input to both schemes is the same, the NN model would require more parameters as well, in order to learn the mapping between the measurements and the location labels. This can be achieved by increasing the model’s complexity, such as by adding more hidden layers or using convolutional layers. Note that, this will directly impact the computational complexity of the NNs.
V-F Simulation study of localization error distribution
@@
Fig. 10b shows the CDF of the location error as estimated using JADE and the algorithm-supervised NN. We observe that the NN performs as good as JADE with sub-meter location accuracy in about 92% of the cases, with slightly better performance for errors less than $50$ cm. For comparison, 98% of the estimates yield an error of <1 m when the NN is trained with true locations. These results are significantly good for two main reasons: because JADE can work with a comparatively large measurement set, and because each measurement location is illuminated by multiple MPCs. Bootstrapping with accurate location labels from JADE thus improves the performance of the NN.
@@
Fig. 10c illustrates the localization error heat map in the H-shaped room, for all test locations, as estimated using the algorithm-supervised NN. Blue hues represent low estimation errors (<1 m), light blue to green hues represent errors between 1 m and 2 m, whereas yellow to red hues convey errors between 3 m and 6 m. The results confirm that the localization error is <1 m throughout the room, except for slightly higher errors in the bottom-right section. A similar situation occurs near the top-right corner of the middle section. The main reason is that the training dataset is not collected at uniformly distributed locations, and may thus be sparser in some areas, which also explains why the heat map is not symmetrical.
@@
In general our performance evaluation shows that a tiny algorithm-supervised NN can accurately localize a user in a complex indoor environment in the presence of two key factors: $i)$ richness of input features, coming either from LoS paths, NLoS paths, or both; and $ii)$ richness of training data, so that the NN observes sufficiently many error-prone samples.
VI Discussion and Conclusions
We presented an algorithm-supervised approach to train a tiny NN to localize a mmWave device using ADoA measurements in an indoor environment. This relieves the burdensome task of collecting training data by harvesting location labels from a bootstrapping localization algorithm. We evaluated the performance of our scheme via channel measurements from a NIST mmWave channel sounder. The channel measurements were processed using the SAGE algorithm to obtain all the MPCs at the receiver. We emphasize that, to the best of our knowledge, this is the first algorithm that processes a full, rich, and complex indoor channel information rather than using information pre-filtered by other algorithms or device firmwares. To identify dominant MPCs reflected off different surfaces, we proposed a recursive clustering algorithm based on DBSCAN. The azimuthal AoAs associated with the cluster centroids were used to compute ADoAs. Our performance evaluation shows that tiny NNs can accurately estimate the location of the receiver when trained with true locations. While the error inevitably increases in the algorithm-supervised case, our scheme still achieves sub-meter error in about 75% of the cases, with a much lower computational complexity than the bootstrapping algorithm. Additional simulation results confirm that the key elements that enable high accuracy are richness of illumination from multiple MPCs and training data.
Our future work will address several challenges. One key challenge is to develop schemes that can automatically decide when to switch between the bootstrapping algorithm JADE and the tiny NN model. It is crucial to run JADE only up to the point where its computational complexity remains within acceptable limits for real-time environments. Automatically determining this optimal point will facilitate seamless transitions between the two schemes during runtime. Additionally, we will focus on developing techniques to dynamically associate AoAs of dominant MPCs with ambient scatterers and anchors in more complex and dynamic indoor environments. Unlike our current indoor setup, dynamic scatterers in complex-shaped rooms can result in a large number of anchors. Correctly associating the ADoAs in such environments will be crucial for both the bootstrapping algorithm and the NN. Addressing these challenges will further improve robustness and accuracy in general indoor scenarios.